03. Two charges 12uC, 8uC are 10 cm apart work done in bringing them 4 cm closer iS

Grade:12

1 Answers

Arun
5 years ago
Dear Prem
q1=12×10^-6,q2=8×10^-6.
Wkt,potential when the are at R1(10 cm)
V1=kq1q2/R1=86.4×10^-1~=86×10^-1.
V2=kq1q2/R2=216×10^-1.
W.D=V2-V1=(216×10^-1)-(86×10^-1)=130×10^-1
                                                                =13J
 
 
Regards
Arun